# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ 3.6 Locating singularities of Painlevé II using transseries
As we see in figures 811, for each solution of (43)–(45) with $\alpha\notin\mathbb{N}$, there is a lattice of simple-pole singularities in the upper half $\xi$ plane. We are interested here in the string of singularities that align themselves in an almost-horizontal direction and tend towards the anti-Stokes line $\mathrm{Im}(\xi)=0$, $\mathrm{Re}(\xi)<0$ ($\mathrm{arg}(\xi)=-\pi$) in the far field. These are important as they will have the strongest influence on the real-line solution $u(x,t)$ of (1) for $x<0$ in the small-time limit. Indeed, our working hypothesis is that these singularities link directly to the important properties of dispersive waves, such as their speed and wavelength. In this subsection, we summarise how to apply a standard transseries expansion to estimate the location of these singularities in the far field. The approach we take relies on transseries and is equivalent to that of Costin & Costin [24], who briefly use a version of PII as an example. Other studies that apply transseries methodology for locating singularities of PII are for the homogeneous version with $\alpha=0$, for which some of the details are different [7, 61, 72].
The starting point is to note that the solution to (43)–(45) has the behaviour
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}+\ldots+\frac{\sigma_{1}}{\xi^{1/4}}\,\mathrm{e}^{2\xi^{3/2}/3}\quad\mbox{as}\quad|\xi|\rightarrow\infty,\quad-\pi<\mathrm{arg}(\xi)<-2\pi/3$ (62)
$\sigma_{1}\neq 0,\quad-\pi<\mathrm{arg}(\xi)<-2\pi/3,$ (63)
where the Stokes constant $\sigma_{1}$ in (62)–(63) is given explicitly by
$\sigma_{1}=\frac{\sin\pi\alpha}{\pi^{1/2}}$ (64)
[38, 51]. Clearly, the exponential term in (62) is exponentially small in the sector $\pi<\mathrm{arg}(\xi)<-2\pi/3$, and can be interpreted as the leading-order behaviour of the non-perturbative part of the asymptotic expansion. (An abbreviated argument for (64) is to see from (42) that the terms in the algebraic terms in the asymptotic series for $F$ behave as
@@
As we explain in appendix C.1, we can extend (62) to a transseries of the form
$F\sim\sum_{n=0}^{\infty}\sigma_{1}^{n}\,\mathrm{e}^{2n\xi^{3/2}/3}\sum_{m=0}^{\infty}\frac{F_{m}^{(n)}}{\xi^{-1/2+3n/4+3m/2}}=\sum_{n=0}^{\infty}\left(\frac{\sigma_{1}\,\mathrm{e}^{2\xi^{3/2}/3}}{\xi^{3/4}}\right)^{n}\sum_{m=0}^{\infty}\frac{F_{m}^{(n)}}{\xi^{-1/2+3m/2}},$ (65)
where $F_{2m}^{(0)}=0$, $F_{0}^{(2n)}=0$, $F_{0}^{(1)}=1$ and, comparing with the notation in (40), $F_{2m+1}^{(0)}=\alpha b_{m}$. The transseries (65) is valid in the sector $\pi<\mathrm{arg}(\xi)<-2\pi/3$, as each term $\mathrm{e}^{2n\xi^{3/2}/3}$ is exponentially smaller than the previous one as $n$ increases.
@@
$\tau=\frac{\sigma_{1}\,\mathrm{e}^{2\xi^{3/2}/3}}{\xi^{3/4}}.$ (66)
The representation (65) remains well-ordered provided $|\tau|\ll 1$, which is certainly true for $|\xi|\gg 1$ in the sector $\pi\leq\mathrm{arg}(\xi)<-2\pi/3$. Thus, in this region we can swap the order of summation, so that
$F\sim\sum_{m=0}^{\infty}\frac{1}{\xi^{-1/2+3m/2}}\sum_{n=0}^{\infty}\tau^{n}F_{m}^{(n)},$ (67)
@@
$F\sim\sum_{m=0}^{\infty}\frac{G_{m}(\tau)}{\xi^{-1/2+3m/2}},$ (68)
where the $G_{m}(\tau)$ are functions of $\tau$ that for $\pi\leq\mathrm{arg}(\xi)<-2\pi/3$ are given by the second (convergent) summation in (67). Crucially, the asymptotic representation (68) applies regardless of whether $|\tau|$ is small, which means it applies as we cross the negative real $\xi$ axis from below (where $|\tau|$ is exponentially small) to above (where $|\tau|$ is exponentially large). In particular, we shall apply (68) in the neighbourhood of the string of singularities that align themselves slightly above the negative real $\xi$ axis for large $|\xi|$.
Thus, near the negative real $\xi$ axis, we substitute (68) into PII (36) to derive odes for $G_{m}$, which are subject to boundary conditions that come from matching into the lower half plane, namely
$G_{m}\sim F_{m}^{(0)}+F_{m}^{(1)}\tau+F_{m}^{(2)}\tau^{2}+\ldots\quad\mbox{as}\quad\tau\rightarrow 0.$ (69)
@@
$F\sim\frac{4\tau\xi^{1/2}}{4-\tau^{2}}=\frac{4\sigma_{1}\xi^{-3/4}\,\mathrm{e}^{2\xi^{3/2}/3}\xi^{1/2}}{4-\sigma_{1}^{2}\xi^{-3/2}\,\mathrm{e}^{4\xi^{3/2}/3}}\quad\mbox{as}\quad|\xi|\rightarrow\infty,$ (70)
@@
$\xi_{0}\sim-(3\pi n)^{2/3}+\frac{1}{(3\pi n)^{1/3}}\left(\frac{3\pi}{4}+\frac{\mathrm{i}}{2}\ln(3\pi n)+\mathrm{i}\log\left(\frac{\tau_{0}^{(0)}}{\sigma_{1}}\right)\right)\quad\mbox{as}\quad n\rightarrow\infty,$ (72)
@@
For our purposes, it is important to note that the Stokes constant $\sigma_{1}$ in (64) is real, but may be positive or negative, depending on $\alpha$. Further, $\sigma_{1}$ vanishes for $\alpha\in\mathbb{N}$. Thus, given our two choices $\tau_{0}^{(0)}=\pm 2$, the term $\tau_{0}^{(0)}/\sigma_{1}$ can also be positive or negative, provided $\alpha\notin\mathbb{N}$. With this in mind, we can rewrite (72) for the two separate cases as
@@
$\xi_{0}\sim-(3\pi n)^{2/3}+\frac{1}{(3\pi n)^{1/3}}\left(\frac{3\pi}{4}+\frac{\mathrm{i}}{2}\ln\left(\frac{12\pi^{2}n}{\sin^{2}\pi\alpha}\right)\right)\quad\mbox{as}\quad n\rightarrow\infty\quad\mbox{for}\quad\tau_{0}^{(0)}/\sigma_{1}>0,$ (73)
@@
$\xi_{0}\sim-(3\pi n)^{2/3}+\frac{1}{(3\pi n)^{1/3}}\left(-\frac{\pi}{4}+\frac{\mathrm{i}}{2}\ln\left(\frac{12\pi^{2}n}{\sin^{2}\pi\alpha}\right)\right)\quad\mbox{as}\quad n\rightarrow\infty\quad\mbox{for}\quad\tau_{0}^{(0)}/\sigma_{1}<0.$ (74)
@@
For example, if we choose $\alpha=1/2$, then $\sigma_{1}>0$. With this parameter choice, poles with $\tau_{0}^{(0)}=2$ (those with residue 1-1) are located using (73), while poles with $\tau_{0}^{(0)}=-2$ (residue +1+1) are located using (74). We show this case in figure 12, where we see the asymptotic formulae (73)–(74) do an excellent job of estimating the pole locations for moderately small values of $n$ even though they hold formally in the limit $n\rightarrow\infty$. (Note that by considering $G_{1}(\tau)$ in (68), we can extend (73)–(74) to include higher-order corrections. See appendix C.3.)
@@
Figure 12: Phase portrait of a numerical solution of (43)–(45), computed for $\alpha=1/2$. The black and white circles denote estimates of poles with residues 1-1 and +1+1 computed via (73) and (74), respectively. The closest two circles to the origin are computed using $n=1$, the next closest pair with $n=2$, and so on.
@@
We make two further observations about our asymptotic formulae (73)–(74). First, the poles alternate between residues +1+1 and 1-1 and, for very large $n$, are separated by an asymptotic distance $\pi^{2/3}/(3n)^{1/3}$. Second, these formulae assume that $\ln|\sigma_{1}|=\mathcal{O}(1)$. For values of $\alpha$ extremely close to a natural number $N$, $\sigma_{1}$ becomes sufficiently small in magnitude that $\ln|\sigma_{1}|$ becomes large and these formulae need reconsidering. Correspondingly, as $\alpha$ approaches $N$, the lattice of poles “moves away” from the origin, leaving behind a block of $N^{2}$ poles that are generated by the relevant rational solution for $\alpha=N$.
@@
It is worth recalling that, in terms of our original KdV problem, double poles of solutions $u(x,t)$, located at $x=s(t)$, have the small-time behaviour (29). Thus, we can use (73) and (74) for $\tau_{0}^{(0)}=-2$, to describe the leading-order location of double poles of $u(x,t)$ that propagate almost horizontally to the left like $s(t)\sim x_{0}-3\pi^{2/3}n^{2/3}t^{1/3}$. Each of these can be associated with a crest of the dispersive waves on the real line, linking real-valued behaviour with poles in the complex plane.
@@ 4 Check of early-time behaviour using $N$-soliton solutions
As a check on our small-time asymptotics, we consider the exact $N$-soliton solutions that comes from the initial condition (4) with $A_{0}=-N(N+1)$. For these examples, there are $N(N+1)/2$ double-pole singularities that emerge from each singularity of the initial condition at $x_{0}=(n+\frac{1}{2})\pi\mathrm{i}$. While initially these move out from each $x_{0}$ in the directions and speeds given by the rational solutions of $f_{0}$ (see (52)–(55), for example), they will ultimately rearrange themselves so that $N$ of them align with the fastest soliton, N1N-1 align with the next fastest, and so on, until only $1$ singularity from each $x_{0}$ is associated with the slowest soliton.
At this point it is worth repeating that we have used a change of variables (31) to rewrite our inner problem for $f_{0}(\xi)$ in terms of solutions $F(\xi)$ of PII. By inverting this transformation, we observe that, of all the simple poles of a solution for $F$, only those with residue +1+1 correspond to double poles of $f_{0}$; the simple poles of $F$ with reside 1-1 do not correspond to singularities of $f_{0}$ and therefore are not associated with singularities of our original KdV problem. For the rational solutions of PII, there are $N^{2}$ simple poles, but only $N(N+1)/2$ of these are double poles of the corresponding rational solutions of $f_{0}$.
@@ 4.1 $1$-soliton solution
An obvious example is a 1-soliton solution $u(x,t)=2\,\mathrm{sech}^{2}(x-4t)$, which evolves from the initial condition $u_{0}=2\,\mathrm{sech}^{2}x$, that is, from (4) with $A_{0}=-2$. The 1-soliton solution is simply a constant-shape travelling wave moving from left to right with speed $4$. Clearly, the solution has poles where $\cosh(x-4t)=0$, which are located at
$s(t)=\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}+4t,$ (75)
@@
With this representation, we see poles of $u$ can be determined via the zeros of $U_{1}$, giving the same result as (75), as expected.
@@
we see from (21) that $A_{0}=-2$, $A_{1}=0$ and $A_{2}=2/3$. Following this very special case in appendix B.4.2, our analysis predicts that the early time behaviour of singularity is described by (127), namely $s(t)\sim x_{0}+6A_{2}t$ as $t\rightarrow 0^{+}$. Given $A_{2}=2/3$, this predicted speed is $6A_{2}=4$, which agrees with the exact solution for the 1-soliton solution. Thus, while the 1-soliton solution itself is trivial in the sense that we can easily determine the small-time behaviour of the complex singularities without any of our asymptotic analysis, this check of (127) provides support for the details in appendix B.4 (which apply for other, more complicated, initial conditions that also have $A_{0}=-2$).

### v2

@@ 3.7 Locating singularities of Painlevé II using transseries
As we see in figures 811, for each solution of (3.23)–(3.25) with $\alpha\notin\mathbb{N}$, there is a lattice of simple-pole singularities in the upper half $\xi$ plane. We are interested here in the string of singularities that align themselves in an almost-horizontal direction and tend towards the anti-Stokes line $\mathrm{Im}(\xi)=0$, $\mathrm{Re}(\xi)<0$ ($\mathrm{arg}(\xi)=-\pi$) in the far field. These are important as they will have the strongest influence on the real-line solution $u(x,t)$ of (1.1) for $x<0$ in the small-time limit. Indeed, our working hypothesis is that these singularities link directly to the important properties of dispersive waves, such as their speed and wavelength. In this subsection, we summarise how to apply a standard transseries expansion to estimate the location of these singularities in the far field. The approach we take relies on transseries and is equivalent to that of Costin & Costin [29], who briefly use a version of PII as an example. Other studies that apply transseries methodology for locating singularities of PII are for the homogeneous version with $\alpha=0$, for which some of the details are different [8, 70, 82].
The starting point is to note that the solution to (3.23)–(3.25) has the behaviour
$F\sim\frac{\alpha}{\xi}+\frac{2\alpha(1-\alpha^{2})}{\xi^{4}}+\ldots+\frac{\sigma_{1}}{\xi^{1/4}}\,\mathrm{e}^{2\xi^{3/2}/3}\quad\mbox{as}\quad|\xi|\rightarrow\infty,\quad-\pi<\mathrm{arg}(\xi)<-2\pi/3$ (3.43)
$\sigma_{1}\neq 0,\quad-\pi<\mathrm{arg}(\xi)<-2\pi/3,$ (3.44)
where the Stokes constant $\sigma_{1}$ in (3.43)–(3.44) is given explicitly by
$\sigma_{1}=\frac{\sin\pi\alpha}{\pi^{1/2}}$ (3.45)
[44, 58]. Clearly, the exponential term in (3.43) is exponentially small in the sector $\pi<\mathrm{arg}(\xi)<-2\pi/3$, and can be interpreted as the leading-order behaviour of the non-perturbative part of the asymptotic expansion. (An abbreviated argument for (3.45) is to see from (3.22) that the terms in the algebraic terms in the asymptotic series for $F$ behave as
@@
As we explain in appendix C.1, we can extend (3.43) to a transseries of the form
$F\sim\sum_{n=0}^{\infty}\sigma_{1}^{n}\,\mathrm{e}^{2n\xi^{3/2}/3}\sum_{m=0}^{\infty}\frac{F_{m}^{(n)}}{\xi^{-1/2+3n/4+3m/2}}=\sum_{n=0}^{\infty}\left(\frac{\sigma_{1}\,\mathrm{e}^{2\xi^{3/2}/3}}{\xi^{3/4}}\right)^{n}\sum_{m=0}^{\infty}\frac{F_{m}^{(n)}}{\xi^{-1/2+3m/2}},$ (3.46)
where $F_{2m}^{(0)}=0$, $F_{0}^{(2n)}=0$, $F_{0}^{(1)}=1$ and, comparing with the notation in (3.20), $F_{2m+1}^{(0)}=\alpha b_{m}$. The transseries (3.46) is valid in the sector $\pi<\mathrm{arg}(\xi)<-2\pi/3$, as each term $\mathrm{e}^{2n\xi^{3/2}/3}$ is exponentially smaller than the previous one as $n$ increases.
@@
$\tau=\frac{\sigma_{1}\,\mathrm{e}^{2\xi^{3/2}/3}}{\xi^{3/4}}.$ (3.47)
The representation (3.46) remains well-ordered provided $|\tau|\ll 1$, which is certainly true for $|\xi|\gg 1$ in the sector $\pi\leq\mathrm{arg}(\xi)<-2\pi/3$. Thus, in this region we can swap the order of summation, so that
$F\sim\sum_{m=0}^{\infty}\frac{1}{\xi^{-1/2+3m/2}}\sum_{n=0}^{\infty}\tau^{n}F_{m}^{(n)},$ (3.48)
@@
$F\sim\sum_{m=0}^{\infty}\frac{G_{m}(\tau)}{\xi^{-1/2+3m/2}},$ (3.49)
where the $G_{m}(\tau)$ are functions of $\tau$ that for $\pi\leq\mathrm{arg}(\xi)<-2\pi/3$ are given by the second (convergent) summation in (3.48). Crucially, the asymptotic representation (3.49) applies regardless of whether $|\tau|$ is small, which means it applies as we cross the negative real $\xi$ axis from below (where $|\tau|$ is exponentially small) to above (where $|\tau|$ is exponentially large). In particular, we shall apply (3.49) in the neighbourhood of the string of singularities that align themselves slightly above the negative real $\xi$ axis for large $|\xi|$.
Thus, near the negative real $\xi$ axis, we substitute (3.49) into PII (3.14) to derive odes for $G_{m}$, which are subject to boundary conditions that come from matching into the lower half plane, namely
$G_{m}\sim F_{m}^{(0)}+F_{m}^{(1)}\tau+F_{m}^{(2)}\tau^{2}+\ldots\quad\mbox{as}\quad\tau\rightarrow 0.$ (3.50)
@@
$F\sim\frac{4\tau\xi^{1/2}}{4-\tau^{2}}=\frac{4\sigma_{1}\xi^{-3/4}\,\mathrm{e}^{2\xi^{3/2}/3}\xi^{1/2}}{4-\sigma_{1}^{2}\xi^{-3/2}\,\mathrm{e}^{4\xi^{3/2}/3}}\quad\mbox{as}\quad|\xi|\rightarrow\infty,$ (3.51)
@@
$\xi_{0}\sim-(3\pi n)^{2/3}+\frac{1}{(3\pi n)^{1/3}}\left(\frac{3\pi}{4}+\frac{\mathrm{i}}{2}\ln(3\pi n)+\mathrm{i}\log\left(\frac{\tau_{0}^{(0)}}{\sigma_{1}}\right)\right)\quad\mbox{as}\quad n\rightarrow\infty,$ (3.53)
@@
For our purposes, it is important to note that the Stokes constant $\sigma_{1}$ in (3.45) is real, but may be positive or negative, depending on $\alpha$. Further, $\sigma_{1}$ vanishes for $\alpha\in\mathbb{N}$. Thus, given our two choices $\tau_{0}^{(0)}=\pm 2$, the term $\tau_{0}^{(0)}/\sigma_{1}$ can also be positive or negative, provided $\alpha\notin\mathbb{N}$. With this in mind, we can rewrite (3.53) for the two separate cases as
@@
$\xi_{0}\sim-(3\pi n)^{2/3}+\frac{1}{(3\pi n)^{1/3}}\left(\frac{3\pi}{4}+\frac{\mathrm{i}}{2}\ln\left(\frac{12\pi^{2}n}{\sin^{2}\pi\alpha}\right)\right)\quad\mbox{as}\quad n\rightarrow\infty\quad\mbox{for}\quad\tau_{0}^{(0)}/\sigma_{1}>0,$ (3.54)
@@
$\xi_{0}\sim-(3\pi n)^{2/3}+\frac{1}{(3\pi n)^{1/3}}\left(-\frac{\pi}{4}+\frac{\mathrm{i}}{2}\ln\left(\frac{12\pi^{2}n}{\sin^{2}\pi\alpha}\right)\right)\quad\mbox{as}\quad n\rightarrow\infty\quad\mbox{for}\quad\tau_{0}^{(0)}/\sigma_{1}<0.$ (3.55)
@@
For example, if we choose $\alpha=1/2$, then $\sigma_{1}>0$. With this parameter choice, poles with $\tau_{0}^{(0)}=2$ (those with residue 1-1) are located using (3.54), while poles with $\tau_{0}^{(0)}=-2$ (residue +1+1) are located using (3.55). We show this case in figure 12, where we see the asymptotic formulae (3.54)–(3.55) do an excellent job of estimating the pole locations for moderately small values of $n$ even though they hold formally in the limit $n\rightarrow\infty$. (Note that by considering $G_{1}(\tau)$ in (3.49), we can extend (3.54)–(3.55) to include higher-order corrections. See appendix C.3.)
@@
Figure 12: Phase portrait of a numerical solution of (3.23)–(3.25), computed for $\alpha=1/2$. The black and white circles denote estimates of poles with residues 1-1 and +1+1 computed via (3.54) and (3.55), respectively. The closest two circles to the origin are computed using $n=1$, the next closest pair with $n=2$, and so on.
@@
We make two further observations about our asymptotic formulae (3.54)–(3.55). First, the poles alternate between residues +1+1 and 1-1 and, for very large $n$, are separated by an asymptotic distance $\pi^{2/3}/(3n)^{1/3}$. Second, these formulae assume that $\ln|\sigma_{1}|=\mathcal{O}(1)$. For values of $\alpha$ extremely close to a natural number $N$, $\sigma_{1}$ becomes sufficiently small in magnitude that $\ln|\sigma_{1}|$ becomes large and these formulae need reconsidering. Correspondingly, as $\alpha$ approaches $N$, the lattice of poles “moves away” from the origin, leaving behind a block of $N^{2}$ poles that are generated by the relevant rational solution for $\alpha=N$.
@@
It is worth recalling that, in terms of our original KdV problem, double poles of solutions $u(x,t)$, located at $x=s(t)$, have the small-time behaviour (3.7). Thus, we can use (3.54) and (3.55) for $\tau_{0}^{(0)}=-2$, to describe the leading-order location of double poles of $u(x,t)$ that propagate almost horizontally to the left like $s(t)\sim x_{0}-3\pi^{2/3}n^{2/3}t^{1/3}$. Each of these can be associated with a crest of the dispersive waves on the real line, linking real-valued behaviour with poles in the complex plane.
@@ 4 Check of early-time behaviour using $N$-soliton solutions
As a check on our small-time asymptotics, we consider the exact $N$-soliton solutions that comes from the initial condition (1.4) with $A_{0}=-N(N+1)$. For these examples, there are $N(N+1)/2$ double-pole singularities that emerge from each singularity of the initial condition at $x_{0}=(n+\frac{1}{2})\pi\mathrm{i}$. While initially these move out from each $x_{0}$ in the directions and speeds given by the rational solutions of $f_{0}$ (see (3.33)–(3.36), for example), they will ultimately rearrange themselves so that $N$ of them align with the fastest soliton, N1N-1 align with the next fastest, and so on, until only $1$ singularity from each $x_{0}$ is associated with the slowest soliton.
At this point it is worth repeating that we have used a change of variables (3.9) to rewrite our inner problem for $f_{0}(\xi)$ in terms of solutions $F(\xi)$ of PII. By inverting this transformation, we observe that, of all the simple poles of a solution for $F$, only those with residue +1+1 correspond to double poles of $f_{0}$; the simple poles of $F$ with reside 1-1 do not correspond to singularities of $f_{0}$ and therefore are not associated with singularities of our original KdV problem. For the rational solutions of PII, there are $N^{2}$ simple poles, but only $N(N+1)/2$ of these are double poles of the corresponding rational solutions of $f_{0}$.
@@ 4.1 $1$-soliton solution
An obvious example is a 1-soliton solution $u(x,t)=2\,\mathrm{sech}^{2}(x-4t)$, which evolves from the initial condition $u_{0}=2\,\mathrm{sech}^{2}x$, that is, from (1.4) with $A_{0}=-2$. The 1-soliton solution is simply a constant-shape travelling wave moving from left to right with speed $4$. Clearly, the solution has poles where $\cosh(x-4t)=0$, which are located at
$s(t)=\left(n+{\textstyle\frac{1}{2}}\right)\pi\mathrm{i}+4t,$ (4.1)
@@
With this representation, we see poles of $u$ can be determined via the zeros of $U_{1}$, giving the same result as (4.1), as expected.
@@
we see from (2.15) that $A_{0}=-2$, $A_{1}=0$ and $A_{2}=2/3$. Following this very special case in appendix B.4.2, our analysis predicts that the early time behaviour of singularity is described by (B.34), namely $s(t)\sim x_{0}+6A_{2}t$ as $t\rightarrow 0^{+}$. Given $A_{2}=2/3$, this predicted speed is $6A_{2}=4$, which agrees with the exact solution for the 1-soliton solution. Thus, while the 1-soliton solution itself is trivial in the sense that we can easily determine the small-time behaviour of the complex singularities without any of our asymptotic analysis, this check of (B.34) provides support for the details in appendix B.4 (which apply for other, more complicated, initial conditions that also have $A_{0}=-2$).